OpenStudy (igreen):

Suppose a segment of mitochondrial DNA (mtDNA) is compared between two similar modern-day species. It is known that this segment of mtDNA changes at a steady rate of one base pair every 4 million years. Compare the sequences of the two mtDNA segments given below. Species A: GCACTAAGCATCGATTT Species B: GCACCAGGCACTGGTTC Use the basic principle of molecular clocks to calculate how long ago the two species diverged from a common ancestor.

OpenStudy (igreen):

Here's a similar question. @Jhannybean

OpenStudy (igreen):

This is what I got: \(6~differences \cdot \dfrac{4~million~years}{1~mutation} = 24~million~years~ago\)

OpenStudy (jhannybean):

Oh this is just doing the reverse process.

OpenStudy (jhannybean):

Is that one of the answer choices?

OpenStudy (igreen):

Yep.

OpenStudy (igreen):

12 million years ago 8 million years ago 48 million years ago 24 million years ago

OpenStudy (jhannybean):

And the others? They probably don't make sense, anyway.

OpenStudy (jhannybean):

So backtracking, last question we found the initial set of differences that was 5, but we needed 8. 5/80 = 1/16 , leaving us with 3 more differences. Within an approximation of 80 years, we found about how many years it would take us to fully complete 8 change of bases.

OpenStudy (jhannybean):

With that we got what would be approx 48 years, right?

OpenStudy (igreen):

Yep.

OpenStudy (jhannybean):

So yes, I think your approach is correct.

OpenStudy (igreen):

Okay, thanks.

OpenStudy (jhannybean):

You're saying that it'll take 24 mil. years to make 6 changes, as compared to the 4 years to make 1 change in the sequence.

OpenStudy (igreen):

Yep. \(Number~of~changes \times rate~of~mutation = how~long~ago~the~two~species\) \(~diverged~from~a~common~ancestor.\)

OpenStudy (jhannybean):

Yeah :)

OpenStudy (igreen):

24 was wrong.. @Jhannybean >.> I have no idea why.